# Ctrl-Room: Controllable Text-to-3D Room Meshes Generation with Layout Constraints

### v3

@@ 3.2 Appearance Generation Stage
Given the layout of an indoor scene, we seek to obtain a proper panorama image to represent its appearance. Instead of incrementally generating multi-view images like [12], we generate the entire panorama at once. We utilize ControlNet [48] to generate a high-fidelity panorama conditioned by the input 3D scene layout. After getting the scene panorama, we recover the depth map using the method [32] to reconstruct a textured mesh through Possion reconstruction [14] and MVS-texture [40].
@@
Fine-tuning ControlNet. ControlNet is a refined Stable Diffusion [25] model conditioned on an extra 2D input. To condition ControlNet on the scene layout, we convert the bounding box representation into a 2D semantic layout panorama through equirectangular projection. In this way, we get a pair of RGB and semantic layout panoramic images for each scene. However, the pre-trained ControlNet-Segmentation [9] is designed for perspective images, and cannot be directly applied to panoramas. Thus, we fine-tune it with our pairwise RGB-Semantic layout panoramas on the Structured3D [49]. As the volume of Structured3D is limited, we apply several augmentation techniques for the training data, including standard left-right flipping, horizontal rotation, and Pano-Stretch [36].
@@
Loop-consistent Sampling. A panorama should be loop-consistent. In other words, its left and right should be seamlessly connected. Although the panoramic horizontal rotation in data augmentation may improve the model’s implicit understanding of the expected loop consistency, it lacks explicit constraints and might still produce inconsistent results. Therefore, we propose an explicit loop-consistent sampling mechanism in the denoising process of the latent diffusion model. As shown in Fig. 3 (b), we rotate both the input layout panorama and the denoised image by $\gamma$ degree in the sampling process, which applies explicit constraints for the loop consistency during denoising. A concurrent work [44] also uses a similar method for panoramic outpainting. More qualitative results in  Fig. 9 verify that our simple loop-consistent sampling method achieves good results without introducing additional learnable parameters.
@@ 3.3 Mask-guided Editing
A user can modify the generated 3D room by changing the position, semantic class, and size of object bounding boxes. Our method will update the panorama accordingly to reflect the user’s edits in 3D space, achieving zero-shot editing without expensive re-training.
The editing should achieve two goals, i.e. altering the content according to the user’s input, and maintaining appearance consistency of the scene objects. We propose a mask-guided image editing as illustrated Fig. 4, where a chair’s position is moved. In the following, we will explain our method with this example. We denote the semantic panorama from the edited scene as $P_{\rm edited}$, then we derive the guidance masks based on its difference from the original one $P_{\rm ori}$. The source mask $\mathbf{m}_{\rm src}$ shows the position of the original chair, and the target mask $\mathbf{m}_{\rm tar}$ indicates the location of the moved chair, and the inpainting mask $\mathbf{m}_{\rm inpaint}=\{m|m\in\mathbf{m}_{\rm src}\;{\rm and}\;m\notin\mathbf{m}_{\rm tar}\}$ is the unoccluded region. Given these guidance masks, our method includes two steps: the inpainting step and the optimization step.
We use $\mathbf{x^{{\rm ori}}_{0}}$ to denote the original image. During the inpainting step, we replace pixels outside the inpainting mask $\mathbf{m}_{\rm inpaint}$ with $\mathbf{x^{{\rm ori}}_{t}}$ and store $\mathbf{m}_{\rm inpaint}$ based on the edited semantic panorama $P_{\rm edited}$. This straightforward approach ensures that the region outside the mask remains unchanged and the area inside the mask is accurately inpainted. We design an optimization step to maintain the appearance of furniture before and after movement and rescaling operations. Recent work DIFT [38], has shown that learned features from the diffusion network enable strong semantic correspondence. Therefore, we ensure consistency between the original and moved furniture by requiring their latent features to be consistent. For more details of the Inpainting Step and Optimization Step, please refer to our supplementary file, Sec.1.2.
@@ 4 Experiments
We evaluate Ctrl-Room on three tasks: layout generation, panoramic image generation, and 3D Room generation. We first describe the experimental settings and then validate our method by comparing it with previous methods quantitatively and qualitatively. We further show various scene editing results to demonstrate the flexible control of our method.
@@ 4.1 Experiment Setup
Dataset: We train and evaluate our method on the 3D indoor scene dataset Structured3D [49], which consists of 3,500 houses with 21,773 rooms designed by professional artists. Photo-realistic 2D renderings, the labeled room layout, and 3D bounding boxes of furniture are provided in each room. We parse these boxes for typical indoor rooms, such as the bedroom, kitchen, living room, study, and bathroom. Then, we follow [42] to generate text prompts describing the scene layout. Please refer to Appendix Sec.1.3 for more details about data preprocessing. The filtered dataset for training and evaluation consists of 4,961 bedrooms, 1,848 kitchens, 3,039 living rooms, 698 studies, and 1500 bathrooms. For each room type, we use $80\%$ of rooms for training and the remaining for testing. Following DiffuScene [37], we further qualitatively evaluate our layout generation on the 813 living rooms and 900 dining rooms of 3D-FRONT dataset [8].
Metrics: Follow previous work [22, 37], Frechet Inception Distance (FID) [11] and Kernel inception distance (KID) [3] are used to measure the plausibility and diversity of 1,000 synthesized scene layouts. We choose FID, CLIP Score (CS) [24], and Inception Score (IS) [28] to measure the image quality of generated panoramas. We also compare the time cost to synthesize an RGB panorama of size $512\times 1024$. To compare the quality of 3D room models, we follow Text2Room [12] to render images of the 3D room model and measure the CLIP Score (CS) and Inception Score (IS). We further conduct a user study and ask $61$ users to score Perceptual Quality (PQ) and 3D Structure Completeness (3DS) of the final room mesh on scores ranging from 1 to 5.
@@ 4.2.1 Qualitative Comparison
Fig. 5 shows some results generated by different methods. The first row shows a textured 3D room model, and the second row shows some perspective renderings from the room model. As we can see, Text2Li-ght [6] cannot generate a reasonable 3D indoor scene. It even fails to ensure the loop consistency of the generated panorama, which leads to distorted geometry and room model. Both MVDiffusion [39] and Text2Room [12] can generate vivid local images as demonstrated by the perspective renderings in the second row. But they fail to capture the more global scale room layout. Similar effects can be seen from the Fig. 1 (a). These two methods often repeat a dominating object, e.g. a bed in the bedroom or fireplace in the living room, multiple times at different places and violate the room layout constraint. In comparison, our method does not suffer from these problems and generates high-quality results. More examples are provided in the Appendix.
Figure 6: Text-conditioned layout generation on Structured3D. Given the text prompt, our method synthesizes a plausible scene layout that matches the description. The generated layout is represented using different colors to indicate various object categories, such as blue for the sofa and brown for the chair.
@@ 4.2.2 Layout Generation
Table 1: Quantitative Comparison of layout generation on 3D-FRONT. Note that DiffuScene-w-SC uses an additional network to learn a Shape Code for each furniture, facilitating the evaluation process to retrieve a more accurate CAD model for each furniture. Nevertheless, our method outperforms others in the common settings, where only the generated semantic class and size of each furniture are used for retrieval.
@@
Fig. 6 verifies that our layout generation results are plausible and can offer reliable 3D room layout constraints for the following appearance generation stage. As shown in  Fig. 6, our text-conditioned layout generation module can synthesize natural and diverse typical indoor scenes on the testing set of Structured3D. The size and spatial location of the furniture are reasonable, and the relative positions between the furniture pieces are accurately recovered. Additional objects not described in the text are automatically generated according to the scene prior. We also perform additional experiments using free-style text prompts generated by GPT-4V [46] (Appendix Sec.7.8).
@@ 4.2.3 Panorama Generation
Text2Light [6] 56.22 21.45 4.198 81.56 - - 2.732 2.747
MVDiffusion [39] 34.76 23.93 3.21 208.5 - - 3.27 3.437
Text2Room [12] - - - $\geq 9,000$ 25.90 2.90 2.487 2.588
Ours 21.02 22.19 3.56 61.1 25.97 3.14 3.89 3.746
Fig. 7 qualitatively evaluates our generated panoramic images. In Fig. 7, the generated panorama is visualized in a panoramic image viewer to facilitate the user to check the global content. The left side of each column is two zoom-in views, and the right side is the fisheye view. Text2Light [6] suffers from serious inconsistency on the borders of the generated panorama. It also shows a lot of unexpected objects in the image. MVDiffusion [39] suffers from repetitive furniture and fails to synthesize reasonable content for the target room type. In contrast, our method obtains a plausible layout and vivid panorama from the given text prompt.
Table 2 provides quantitative evaluations. We follow MVDiffusion [39] to crop perspective images from the generated panoramas on the test split and evaluate the FID, CS, and IS scores on the cropped multi-view images. In the left part of Table 2, our method achieves the best score in FID, which indicates that our method can better capture the room appearance because of its faithful recovery of the room layout. However, our score on CS is slightly lower than MVDiffusion, which seems insensitive to the number of objects and cannot reflect the quality of room layouts. The IS score depends on the semantic diversity of the cropped images as captured by an image classifier. Text2Light has the best IS score, since its generated indoor scenes often contain unexpected objects.
Fig. 9 shows the performance of our panorama generation module with and without loop-consistent sampling mechanism, indicating the loop-consistent sampling helps the generated panorama obtain better texture consistency.
In terms of running time, our method takes the shortest time. On average, our method takes only 61 seconds to generate a panorama, and another 20 seconds to generate the textured 3D mesh. In comparison, MVDiffusion takes 208 seconds, about 3.5 times longer, to synthesize a panorama. Text2Room needs at least 150 minutes to finish a textured 3D room generation.
@@
We then compare the 3D room models in terms of their rendered images. Because of the expensive running time of Text2Room [12], we only test on 12 examples for this comparison. In this comparison, we further skip Text2light and MVDiffusion since we have compared them on panoramas. As the room layout is better captured with a large FOV, we render 60 perspective images of each scene with a $140^{\circ}$ FOV and evaluate their CS and IS scores respectively. Please refer to Appendix Sec.1.4.1 for more details. The results of this comparison are shown in the middle of Table 2. Our method obtains better scores on both metrics than Text2Room.
We further evaluate the quality of the textured 3D mesh model by user studies. For those panorama generation methods, we utilize the depth estimation work [32] to reconstruct a textured 3D room mesh. Further implementation details can be found in the Appendix Sec.1.4. The results of the user study are shown on the right of Table 2. Users prefer our method over others, for its clear room layout structure and furniture arrangement.
@@ 4.3 Interactive Scene Editing
We demonstrate the scene editing capability of our method in  Fig. 8. In this case, we resize the TV and replace the chair in the generated results. Fig. 1 (b) shows two additional examples of replacing the TV and TV stand and moving the picture upwards. Our method can keep the visual appearance of the moved/resized objects unchanged after editing. More examples can be found in the appendix.
@@ 5 Conclusion
We present Ctrl-Room, a flexible method to achieve structurally plausible and editable 3D indoor scene generation. It consists of two stages, the layout generation stage and the appearance generation stage. In the layout generation stage, we design a scene code to parameterize the scene layout and learn a text-conditioned diffusion model for text-driven layout generation. In the appearance generation stage, we fine-tune a ControlNet model to generate a vivid panorama image of the room with the guidance of the layout. Finally, a high-quality 3D room with a structurally plausible layout and realistic textures can be generated. We conduct extensive experiments to demonstrate that Ctrl-Room outperforms existing methods for 3D indoor scene generation both qualitatively and quantitatively, and supports interactive 3D scene editing.
@@ 6 Limitation
There are still some limitations of Ctrl-Room. Firstly, we only support single-room generation, thus we cannot produce large-scale indoor scenes with multiple rooms. Secondly, the generated 3D room still contains incomplete structures in invisible areas because of the occlusion and poor performance of the panoramic depth estimator. A promising direction is to learn a text-driven diffusion model to produce one or more RGB-D panorama images under the scene layout constraints. Lastly, as we explore injecting 3D scene information into pretrained 2D models, thus we rely on 3D labeled scene dataset to drive the learning and fine-tuning process. Leveraging scene datasets with only 2D labels to learn 3D priors is also a promising direction. We leave the aforementioned limitations for our future efforts.
7 Supplymentary File
In the supplementary file, we first present more details about our scene code diffusion model in Sec. 8, then we elaborate the mask-guided editing method in Sec. 9. Next, we provide our dataset pre-processing, text prompt generation, and implementation details in Sec. 10 and Sec. 11 respectively. Additional experiment results are also illustrated, including panorama generation comparisons in Sec. 12, room mesh comparisons in Sec. 13 and user studies in Sec. 14. Furthermore, we demonstrate that our scene code diffusion model can be trained with free-style text prompts in Sec. 15.

### v4

@@ 3.2 Appearance Generation Stage
Given an indoor scene layout, we seek to generate the 3D textured room model. We achieve this goal by generating panoramic images and reconstructing a panoramic NeRF (PeRF) model from these panoramas. During the panorama generation, instead of incrementally generating multi-view images like [12], we generate the entire panorama at once. We utilize ControlNet [47] to generate a high-fidelity panorama conditioned by the 3D scene layout.
3.2.1 Layout-guided Panorama Generation
@@
Fine-tuning ControlNet. ControlNet controls the image generation of Stable Diffusion [24] model by an extra 2D input. To condition ControlNet on the scene layout, we convert the bounding box representation into a 2D semantic layout panorama through equirectangular projection. In this way, we get a pair of RGB and semantic layout panoramic images for each scene. However, the pre-trained ControlNet-Segmentation [9] is designed for perspective images, and cannot be directly applied to panoramas. Thus, we fine-tune it with our pairwise RGB-Semantic layout panoramas on the Structured3D [48]. As the volume of Structured3D is limited, we apply several augmentation techniques for the training data, including standard left-right flipping, horizontal rotation, and Pano-Stretch [33].
@@
Loop-consistent Sampling. A panorama should be loop-consistent. In other words, its left and right should be seamlessly connected. Although the horizontal rotation in data augmentation may improve the model’s implicit understanding of the expected loop consistency, it lacks explicit constraints and might still produce inconsistent results. Therefore, we propose an explicit loop-consistent sampling mechanism in the denoising process of the latent diffusion model. As shown in Fig. 3 (b), we rotate both the input layout panorama and the denoised image by $\gamma$ degree in the sampling process, which applies explicit constraints for the loop consistency during denoising. A concurrent work [42] also uses a similar method for panoramic outpainting. More qualitative results in supplementary Fig.8 and Fig.9 verify that our simple loop-consistent sampling method achieves good results without introducing additional learnable parameters.
Figure 4: The Layout-guided PeRF takes the input panorama, aligned depth map, and normal map as initialization. Then, a progressive inpainting module is introduced to generate consistent panoramic images at sampled novel views. The progressive inpainting module consists of the layout-guided panorama inpainting and the layout-guided depth estimation module. The final RGB-D panoramic pairs are included as training views to finetune PeRF [38].
3.2.2 Layout-guided PeRF Generation
Since a single panorama is only a partial observation of a scene up to occlusions, lifting a single view into a 3D room becomes complex. Fortunately, our generated layout provides valuable geometric and semantic information to lift the 2D panorama into a 3D model. We propose the layout-guided PeRF, which upgrades the generated panorama aforementioned to a 3D panoramic NeRF [38], enabling multi-view consistent panorama generations guided by the scene layout. Specifically, we start with the layout-guided depth estimation, which recovers the depth map using method [45] and then aligns it to the 3D scene layout leveraging its geometric information. This step corrects the biased depth prediction in the background (wall, ceiling, floor) and preserves objects’ shape in the foreground.
Then, we fit our layout-guided PeRF as illustrated in Fig. 4. Specifically, we initialize the scene NeRF with the panorama $I^{0}$, the aligned depth map $D^{*}$, and the normal map $N^{*}$. We sample new viewpoints in the occupancy grid that do not conflict with the initial furniture arrangement. At the i-th novel view, we render semantic map $S^{i}_{l}$, depth map $D^{i}_{l}$, and instance map $M^{i}_{l}$ from the scene layout, these are then combined with the panoramic rendering $I^{i}_{r}$ and inpainting mask $\mathbf{m}_{\rm inpaint}$ obtained from the NeRF and fed to the layout-guided panorama inpainting module to generate the novel view panorama. Using our fine-tuned ControlNet, it achieves training-free panoramic inpainting, which replaces pixels outside the inpainting mask $\mathbf{m}_{\rm inpaint}$ with $I^{i}_{r}$ and fill $\mathbf{m}_{\rm inpaint}$ based on the semantic map $S^{i}_{l}$. Subsequently, after generating the novel view image, we apply the layout-guided depth estimation and include it as training views for PeRF following their framework [38]. More details and results can be found in the appendix Sec.2.
@@ 3.3 Mask-guided Editing
A user can modify the generated 3D room by changing the position, semantic class, and size of object bounding boxes. The editing should achieve two goals, i.e. altering the content according to the user’s input, and maintaining appearance consistency of the scene objects. We propose a mask-guided image editing, including inpainting step and optimization step as illustrated Fig.6 in supplementary file. The inpainting step fills in the modified area while preserving the rest of the panoramic image. The optimization step focuses on keeping the furniture’s appearance unchanged before and after movement and scaling operations.
We explain our method by taking the example in Fig.6 in supplementary file, where a chair’s position is moved. We denote the semantic panorama from the edited scene as $S_{\rm edited}$, then we derive the guidance masks based on its difference from the original one $S_{\rm ori}$. The source mask $\mathbf{m}_{\rm src}$ shows the position of the original chair, and the target mask $\mathbf{m}_{\rm tar}$ indicates the location of the moved chair, and the inpainting mask $\mathbf{m}_{\rm inpaint}=\{m|m\in\mathbf{m}_{\rm src}\;{\rm and}\;m\notin\mathbf{m}_{\rm tar}\}$ is the unoccluded region. We use $\mathbf{x^{{\rm ori}}_{0}}$ to denote the original image. During the inpainting step, we replace pixels outside the inpainting mask $\mathbf{m}_{\rm inpaint}$ with $\mathbf{x^{{\rm ori}}_{t}}$ and store $\mathbf{m}_{\rm inpaint}$ based on the edited semantic panorama $S_{\rm edited}$. This straightforward approach ensures that the region outside the mask remains unchanged and the area inside the mask is accurately inpainted. In the optimization step, drawing inspiration from DIFT [35], which has shown that learned features from the diffusion network enable strong semantic correspondence, we ensure consistency between the original and moved furniture by requiring their latent features to be consistent. For more details of the Inpainting and Optimization Step, please refer to our supplementary file Sec.3.
@@ 4 Experiments
We evaluate Ctrl-Room on three tasks: layout generation, panorama generation, and 3D Room generation. For those panorama generation methods [6, 36], we recover its depth map using method [45] to reconstruct a textured mesh through Possion reconstruction [14] and MVS-texture [37]. We first describe the experimental settings and then validate our method by comparing it with previous methods quantitatively and qualitatively. We further show various scene editing results to demonstrate the flexible control of our method.
@@ 4.1 Experiment Setup
Dataset: We train and evaluate our method on the 3D indoor scene dataset Structured3D [48], which consists of 3,500 houses with 21,773 rooms designed by professional artists. A single panoramic image and 3D scene layout are provided in each room. We parse the scene layout using oriented bounding boxes for common indoor room types like the bedroom, kitchen, living room, study, and bathroom. Then, we follow [40] to generate text prompts describing the scene layout. The filtered dataset for training and evaluation consists of 4,961 bedrooms, 1,848 kitchens, 3,039 living rooms, 698 studies, and 1500 bathrooms. For each room type, we use $80\%$ of rooms for training and the remaining for testing. Following DiffuScene [34], we further qualitatively evaluate our layout generation on 3D-FRONT dataset [7].
Metrics: Follow previous work [21, 34], Frechet Inception Distance (FID) [11] and Kernel inception distance (KID) [3] are used to measure the plausibility and diversity of 1,000 synthesized scene layouts. We choose FID, CLIP Score (CS) [23], and Inception Score (IS) [27] to measure the image quality of generated panoramas. To compare the quality of 3D room models, we follow Text2Room [12] to render images of the 3D room model and measure the CLIP Score (CS) and Inception Score (IS). We also conduct a user study and ask $61$ users to score Perceptual Quality (PQ) and 3D Structure Completeness (3DS) of the final room mesh on scores ranging from 1 to 5.
More details about data preprocessing, experimental settings, and baseline implementations can be found in supplementary file Sec.4 and Sec.5.
@@ 4.2.1 Qualitative Comparison
Fig. 5 shows some results generated by different methods. The first row shows a textured 3D room model, and the second row shows some perspective renderings from the room model. As we can see, Text2Light [6] fails to ensure the loop consistency of the generated panorama, which leads to distorted geometry and unreasonable room model. Both MVDiffusion [36] and Text2Room [12] can generate vivid local images as demonstrated by the perspective renderings in the second row. But they fail to capture the global room layout. These two methods often repeat a dominating object, e.g. the cabinet in the bedroom appears multiple times at different places and violate the room layout constraint. In comparison, our method does not suffer from these problems and generates high-quality results. More examples are provided in the Fig.12 in supplementary file.
Figure 6: Text-conditioned layout generation on Structured3D. Given the text prompt, our method synthesizes a plausible scene layout that matches the description. The generated layout is represented using different colors to indicate various object categories, such as blue for the sofa and brown for the chair. More results and semantic labels are provided in Fig.10 in supplementary file.
@@ 4.2.2 Layout Generation
Table 1: Quantitative Comparison of layout generation on 3D-FRONT. Note that DiffuScene-w-SC uses an additional network to learn a Shape Code for each furniture, facilitating the evaluation process to retrieve a more accurate CAD model for each furniture. Nevertheless, our method outperforms others in the common settings, where only the generated semantic class and size are used for retrieval.
@@
Fig. 6 verifies that our layout generation results are plausible and can offer reliable 3D scene layout constraints for the following appearance generation stage. As shown in  Fig. 6, our text-conditioned layout generation module can synthesize natural and diverse typical indoor scenes. The size and spatial location of the furniture are reasonable, and the relative positions between the furniture pieces are accurately recovered. Additional objects not described in the text are automatically generated according to the scene prior.
@@ 4.2.3 Panorama Generation
Fig. 7 qualitatively evaluates our generated panoramic images, the image is visualized in a panoramic image viewer to facilitate the user to check the global content. The left side of each column is two zoom-in views, and the right side is the fisheye view. Text2Light [6] suffers from serious inconsistency on the borders of the generated panorama. It also shows a lot of unexpected objects in the image. MVDiffusion [36] suffers from repetitive furniture and fails to synthesize reasonable content for the target room type. In contrast, our method obtains a plausible layout and vivid panorama from the given text prompt.
@@
Text2Light [6] 56.22 21.45 4.198 - - 2.732 2.747
MVDiffusion [36] 34.76 23.93 3.21 - - 3.27 3.437
Text2Room [12] - - - 25.90 2.90 2.487 2.588
Ours 21.02 22.19 3.56 25.97 3.14 3.89 3.746
Table 2 provides quantitative evaluations. We follow MVDiffusion [36] to crop perspective images from the generated panoramas on the test split and evaluate the FID, CS, and IS scores on the cropped multi-view images. In the left part of Table 2, our method achieves the best score in FID, which indicates that our method can better capture the room appearance because of its faithful recovery of the room layout. However, our score on CS is slightly lower than MVDiffusion, which seems insensitive to the number of objects and cannot reflect the room layout quality. The IS score depends on the semantic diversity of the cropped images as captured by an image classifier. Text2Light has the best IS score, since the generations contain unexpected objects.
In Fig.8 of the supplementary file, we also study the performance of our panorama generation module with and without loop-consistent sampling mechanism, the ablation indicates the loop-consistent sampling helps the generated panorama obtain better texture consistency.
@@
We then compare the 3D room models in terms of their rendered images. Because of the expensive running time of Text2Room [12], we only test on 12 examples for this comparison. In this comparison, we further skip Text2light and MVDiffusion since we have compared them on panoramas. As the room layout is better captured with a large FOV, we render 60 perspective images of each scene with a $140^{\circ}$ FOV and evaluate their CS and IS scores respectively. The results of this comparison are shown in the middle of Table 2. Our method obtains better scores on both metrics than Text2Room.
We further evaluate the quality of the textured 3D mesh model by user studies. The results of the user study are shown on the right of Table 2. Users prefer our method over others, for its clear room layout structure and furniture arrangement.
@@ 4.3 Interactive Scene Editing
We demonstrate the scene editing capability of our method in  Fig. 8. In this case, we resize the TV and replace the chair in the generated results. Fig. 1 (b) shows examples of replacing the TV and TV stand. Our method can keep the visual appearance of the moved/resized objects unchanged after editing. More examples can be found in the appendix.
@@ 5 Conclusion
We present Ctrl-Room, a flexible method to achieve structurally plausible and editable 3D indoor scene generation. It consists of two stages, the layout generation stage and the appearance generation stage. In the layout generation stage, we design a scene code to parameterize the scene layout and learn a text-conditioned diffusion model for text-driven layout generation. In the appearance generation stage, we fine-tune a ControlNet model to generate a vivid panorama image of the room with the guidance of the layout. Finally, a high-quality 3D room with a structurally plausible layout and realistic textures can be generated via the layout-guided panoramic NeRF. We conduct extensive experiments to demonstrate that Ctrl-Room outperforms existing methods for 3D indoor scene generation both qualitatively and quantitatively, and supports interactive 3D scene editing.
@@ 6 Limitation
There are still some limitations of Ctrl-Room. Firstly, we only support single-room generation, thus we cannot produce large-scale indoor scenes with multiple rooms. A promising direction is to learn a text-driven diffusion model to produce more consistent RGB-D panorama images cross multiple rooms under the scene layout constraints. Secondly, as we explore injecting 3D scene information into pretrained 2D models, thus we rely on 3D labeled scene dataset to drive the learning and fine-tuning process. Leveraging scene datasets with only 2D labels to learn 3D priors is also a promising direction. Thirdly, the generated 3D model still contains artifacts and incomplete structures in invisible areas because of the occlusion and poor performance of the panoramic depth estimator. We leave the aforementioned limitations for our future efforts.
[…]
Supplementary Material
In the supplementary file, we first present more details about our scene code diffusion model in Sec. 7, then we elaborate the layout-guided PeRF module and mask-guided editing method in Sec. 8 and Sec. 9, respectively. Next, we provide our dataset pre-processing, text prompt generation, and implementation details in Sec. 10 and Sec. 11 respectively. Additional experiment results are also illustrated, including panorama generation comparisons in Sec. 12, room layout generations and room mesh comparisons in Sec. 13 and user studies in Sec. 14. Furthermore, we demonstrate that our scene code diffusion model can be trained with free-style text prompts in Sec. 15.